OpenStudy (anonymous):

What is the slope of the equation 2y = 6x - 8? y = -3 y = -3x - 2 y = -2x - 3 y = -2 @iGreen can you help with this one as well?

OpenStudy (igreen):

Okay, we want to solve for 'y', so divide 2 to both sides. y = 6/2x - 8/2 What's 6 / 2?

OpenStudy (amorfide):

you need to rearrange the equation into the form y=mx+c where m is the gradient and c is the y intercept

OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, -2) and has a slope of -3? y = -3 y = -3x - 2 y = -2x - 3 y = -2 This was supposed to be the question. I posted the wrong one

OpenStudy (amorfide):

it has a slope of -3 so m=-3 the equation of a line is y=mx+c you know the coefficient of x is -3 there is only one option that fits this

OpenStudy (igreen):

Okay, plug it in point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y + 2 = 3(x - 0)\) Distribute 3 into the parenthessis: \(y + 2 = 3x - 0\) Subtract 2 to both sides, what's 0 - 2?

OpenStudy (igreen):

Oh, my bad that should be: y + 2 = -3x - 0

OpenStudy (anonymous):

So the answer is B

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

Thanks!